john's age is 3 less than three times marc's age the sum of their ages is 57 write a system of equation that relate their age

Answer by ikleyn(52847)

.
This system is

J = 3*M - 3,     (1)
J + M = 57.      (2)


Or, if you want to write it in the standard form,

J - 3M = -3,
J  + M = 57.